OpenStudy (anonymous):

Grace owns a car with a 17-gallon gas tank. after filling it, she used 10 1/3 driving to her mother's. find the number of gallons left. exacts: grace's tank had _ gallons of gas left

OpenStudy (chriss):

convert both numbers to improper fractions multiply 17*3 and put that result over 3 and multiply 10*3 and add the 1 to convert 10 1/3 \[\frac{51}{3}-\frac{31}{3}\] 51-31=20 so your answer is \[\frac{20}{3} = 6\frac{2}{3}\]